OpenStudy (anonymous):

@rational decide if you wanna play this game or not . I toss a coin and you call if you win I give you 200$ if you lose you give me 100$ If I toss 10 times

OpenStudy (rational):

il play if you let me give you the coin

OpenStudy (anonymous):

haha are u sure ?

OpenStudy (rational):

just want to make sure you're not using a custom made coin with some kind of robotic sensor lol

OpenStudy (rational):

definitely, the expected value is positive here, so it is always wise to play this game

OpenStudy (anonymous):

:P ur changing ur mind so quick

OpenStudy (rational):

nope, i said il play the game if you let me give you the coin

OpenStudy (anonymous):

oh so u wont let me toss it ?

OpenStudy (rational):

i want the coin to be unbiased

OpenStudy (anonymous):

haha ok so u would give me 200$ in case i guessed right ?

OpenStudy (rational):

ugh no

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

haha

OpenStudy (rational):

success probability = 1/2 failure probability = 1/2 since you're giving me 200 for success, and i just need to give you 100 for failure, il be making profit in the long therm

OpenStudy (zzr0ck3r):

He wants it to be a fair coin, he is not trying to trade you your fair coin for his unfair coin.

OpenStudy (rational):

exactly ^

OpenStudy (anonymous):

ok @rational i wont play it with u :P at least i would loose 100$

OpenStudy (rational):

in 10 tosses, you will be loosing $500 i guess

OpenStudy (anonymous):

i said at least 100$

OpenStudy (anonymous):

smallest case me to guess wrong 7 times

OpenStudy (anonymous):

worst case to guess wrong 10 times i would lose 1000$

OpenStudy (rational):

if it is your lucky day, you might win all 10 tosses! who knows!

OpenStudy (anonymous):

im lucky in those games :P

OpenStudy (fibonaccichick666):

unless you are capable of flipping with identical spins of the coin...

OpenStudy (fibonaccichick666):

then you may have an issue ;P

OpenStudy (anonymous):

hahaha if i flip the coin i can know what it is without looking xD

OpenStudy (fibonaccichick666):

yep. I know the trick too :)

OpenStudy (anonymous):

hehe

OpenStudy (anonymous):

i use number theory in that trick (even odd ) case xD

OpenStudy (rational):

you want to find probability of making profit ?

OpenStudy (anonymous):

sure

OpenStudy (rational):

this is also a binomial distribution

OpenStudy (anonymous):

yeah i would like to learn it

OpenStudy (rational):

I need to win at least 4 tosses to make profit, yes ?

OpenStudy (anonymous):

yes

OpenStudy (rational):

By winning 4 tosses, il be getting 4*200 = 800 after that even if loose all remaining tosses, il be only giving you 6*100 = 600 il be making 200 profit

OpenStudy (rational):

so simply find the probability of P(X >= 4)

OpenStudy (anonymous):

there is one thing good about tossing coin its black and white property lol like p(win)=p(lose) (for first time ) so its always easy to calculate

OpenStudy (rational):

the probability of making profit for me is 53/64 which is pretty huge !

OpenStudy (anonymous):

yes of losing it was 11/64

OpenStudy (anonymous):

dou know that binomial distribution is derived from beta function ?

OpenStudy (rational):

never took statistics seriously im familiar wid beta function, so it does look interesting... teach me how to derive it

OpenStudy (rational):

\[B(m,n) = \int\limits_0^1 x^{m-1}(1-x)^{n-1}~dx\]

OpenStudy (rational):

Ahh i see.. this gives the area under binomial distribution curve ?

OpenStudy (anonymous):

yes now chance limits of integral to 0-(1-p) and see what u would got

OpenStudy (anonymous):

almost yes

OpenStudy (anonymous):

it gives the whole area

OpenStudy (rational):

step by step please

OpenStudy (rational):

do we need to treat "x" as probability of success and "1-x" as probability of failure ?